# SkinGrip: An Adaptive Soft Robotic Manipulator with Capacitive Sensing for Whole-Limb Bathing Assistance

### v1

@@ Abstract
Robotics presents a promising opportunity for enhancing bathing assistance, potentially to alleviate labor shortages and reduce care costs, while offering consistent and gentle care for individuals with physical disabilities. However, ensuring flexible and efficient cleaning of the human body poses challenges as it involves direct physical contact between the human and the robot, and necessitates simple, safe, and effective control. In this paper, we introduce a soft, expandable robotic manipulator with embedded capacitive proximity sensing arrays, designed for safe and efficient bathing assistance. We conduct a thorough evaluation of our soft manipulator, comparing it with a baseline rigid end effector in a human study involving $12$ participants across $96$ bathing trails. Our soft manipulator achieves an an average cleaning effectiveness of $88.8\%$ on arms and $81.4\%$ on legs, far exceeding the performance of the baseline. Participant feedback further validates the manipulator’s ability to maintain safety, comfort, and thorough cleaning. Video demonstrations are available on our project website  [^1] https://sites.google.com/view/softbathing.
@@ I Introduction
Prior literature has introduced a number of robotic devices to assist a person in bathing their body. The majority of prior work introduces rigid flat bathing tools to contact and clean small surface regions of human limbs [7, 8, 9]. However, these devices present certain challenges: the use of rigid bathing tools that come in contact with human skin may pose safety risks during interactions with individuals who require assistance, and achieving effective cleaning of a limb’s entire cylindrical surface often necessitates large and complex movements with a rigid end effector. Soft deformable devices present a promising alternative in assistive robotics, attributed to their lightweight, safe touch, and high adaptability [10, 11].
@@
In this work, we develop a robotic system to improve safety in human-robot interactions, specifically designed to aid in bathing tasks, as shown in Fig. 1. This system includes a novel soft robotic manipulator—SkinGrip, equipped with two tendon-driven soft fingers. The two fingers are uniquely structured to gently encircle human limbs. Additionally, we present a capacitive servoing control scheme. This approach empowers the SkinGrip to facilitate safe and continuous contact with human limbs, when used with a Stretch RE1 mobile manipulator. Finally, we conduct a robotic bathing human study experiment to evaluate our SkinGrip against a baseline end effector, evaluating cleaning performance and participant perspectives, across participants with varying sized limbs. The study results demonstrate that the proposed manipulator and control strategy were able to efficiently clean the entire surface of a human limb with an average cleaning effectiveness of $88.8\%$ on arms and $81.4\%$ on legs, surpassing the baseline in both cleaning efficiency and user experience. In this work our contributions include:
@@
    We propose a control algorithm for the SkinGrip that ensures effective skin bathing performance and safe physical interaction. This control method integrates multidimensional capacitive sensing to maintain dynamic, full contact with the skin.
@@ II-A Robot-Assisted Bed Bathing
Bed bathing is a crucial aspect of nursing care, and various robot-assisted approaches have been proposed to enhance this process. Zlatintsi and Dometios et al. [12, 13, 14, 15, 16, 17] developed a vision-based system with a soft arm for washing or wiping the back region of a person. King et al. [8] showcased an end effector designed for limb cleaning, operated via a point cloud interface for area selection. Erickson et al. [7, 9] utilized capacitive sensing to track human motion, enabling a mobile manipulator to move a rigid end effector along a human limb. Liu et al. [18] introduced a mesoscale wearable robot that cleans human limbs by navigating over them. Recently, Madan et al. [19] presented RABBIT, a unique robot-assisted bed bathing system that employs multimodal perception and dual compliance for a robotic manipulator to navigate a flat bathing tool along a human limb. Gu et al. [20] introduced a visuo-tactile Transformer-based imitation learning method (VTTB) for bathing assistance, employing multimodal sensing with a Baxter robot to track body contours and perform precise actions across different subjects. Building on these advancements, our work introduces a soft, expandable robotic manipulator that leverages capacitive sensing control for safe and efficient bathing assistance.
@@ II-B Soft Manipulators for Healthcare
Soft robotics has become increasingly prevalent in healthcare and biomedical fields, especially for rehabilitation and assistance. A comprehensive review [21] examines soft robotic hands designed for rehabilitation, revealing that $34\%$ of the $44$ unique devices analyzed are tendon-driven. Polygerinos et al. [22] developed an electromyography (EMG)-activated glove to aid in grasping objects during ADLs. Cianchetti et al. [23] provide a comprehensive overview of soft robotics in biomedical applications, highlighting compliance, safety, and effectiveness in gentle interactions with human limbs, especially under high contact forces. Furthermore, advancements in soft robotics have extended to assistive devices for bathing. Zlatintsi et al. [14] introduced a fluid and tendon-operated soft robotic arm within the I-Support bathing system. Huang et al. [24] presented a soft pneumatic robotic tool for wiping and bathing that can adapt to slight body contours, but can only make contact with small regions of skin at a time.
@@ II-C Capacitive Servoing and Control
Capacitive sensing is widely utilized in healthcare for applications ranging from classifying gait phases with shoe insole sensors [25] to capturing clothing state as a person is getting dressed using capacitive sensors embedded in the garment [26]. Similarly, Erickson et al. [7, 9] developed a capacitive servoing method using an array of capacitive sensors on a flat rigid tool for robotic bathing and dressing assistance. This approach enabled a robot end effector to traverse the contours of a human limb, but was only demonstrated to clean off the top surface of a limb. Building on these concepts, our work also employs capacitive proximity sensing, but we focus on defining control trajectories for a new manipulator designed with soft fingers that can safely encircle and clean the entire circumference of a limb, addressing the limitations of previous designs.
@@ III Design, Fabrication, and Capacitive Sensing
Figure 2: 3D model and prototype of the proposed SkinGrip, as well as the baseline end effector. (a) The SkinGrip is mounted on a Stretch RE1 robot for cleaning tasks. (b) 3D model of the SkinGrip. (c) Real-world of the SkinGrip, equipped with eight copper foils labeled as: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, $6$-right bottom, $7$-left inner, and $8$-right inner. (d) & (e) The SkinGrip equipped with a sponge and then wrapped with a washcloth. (f) 3D model of the baseline end effector. (g) Real-world of the baseline end effector, equipped with six copper foils labeled as: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, and $6$-right bottom. (h) & (i) The baseline end effector equipped with a sponge and then wrapped with a washcloth.
III-A 3D Model and Prototype of the SkinGrip
@@
We contrast the SkinGrip to a baseline rigid flat end effector based on prior work [9, 7], shown in Fig. 2(f)-(i). This baseline end effector consists of six capacitive electrodes arranged in a $2\times 3$ grid, adhered to the bottom surface of the rigid tool, as shown in Fig. 2(g). These copper foils are connected to a Teensy LC microcontroller board. The flat tool is also covered with shape memory foam and a washcloth, shown in Fig. 2(h)-(i). This rigid flat tool was manufactured using a Raise3D E2 printer with PLA material. Finally, we mount the two end effectors to a Stretch RE1 mobile manipulator to perform body bathing along the human limbs, as shown in Fig. 2(a).
III-B End Effector Control from Capacitive Sensing
Our goal is to instruct the robot to maneuver both the SkinGrip and the baseline end effector along a human limb to provide bathing assistance. The Stretch RE1 must control the end effector’s position and orientation relative to a person’s limb, using capacitance measurements obtained from the integrated capacitive proximity sensors.
@@ III-B End Effector Control from Capacitive Sensing
Soft Finger Control In order to define a control strategy for the SkinGrip, we first quantify the relationship between capacitance measurements from the embedded electrodes and displacement of the fingers around a human limb. We have the two fingers of the SkinGrip wrap around an individual’s limbs, which remain stationary throughout the entire procedure. Fig. 3(a) and (b) depict the correlation between capacitance values and tendon line displacement as soft fingers wrap around a human arm and lower leg, respectively. The results show that capacitance values increase with the tendon’s displacement, indicating an increasing level of tightness between the soft finger and the limb. “Full contact” refers to the condition in which the inner surface of the soft finger is in complete contact with the human skin. Furthermore, as the servos continue to apply tension on the tendon, the soft finger compresses the human skin until it achieves “Tight contact”, at which point the capacitance values stabilize. Notably, capacitance measurements from all sensors routinely increase as the fingers continue to tighten around a human limb, and this trend is consistent for both arms and legs.
@@
The robot captures capacitance measurements from each electrode at a frequency of 10 Hz. At each time step, the robot will take control actions based on a windowed observation of the most recent capacitance measurements $\bm{C}^{s}_{t-4:t}\in\mathbb{R}^{8\times 5}$ from the capacitive sensor from the time step $t-4$ to the current time step $t$. Subsequently, we normalize all real-time capacitance values to the range $[0,1]$ using the formula $\bm{C}_{norm}^{s}=\frac{\sum\bm{C}^{s}_{t-4:t}/5-\bm{C}_{min}^{s}}{\bm{C}_{max}^{s}-\bm{C}_{min}^{s}}$ where the superscript $s$ represents the SkinGrip. Normalization of these values, determined by the maximum $\bm{C}_{max}^{s}$ and minimum $\bm{C}_{min}^{s}\in\mathbb{R}^{8}$ capacitance values, ensures consistent sensitivity across different conditions. We establish a threshold $\bm{C}_{th}^{s1}\in\mathbb{R}^{6}$ for six of the eight electrodes, specifically: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, and $6$-right bottom (Fig. 2(c)), to ensure “full contact” between the soft fingers and the human skin. Here, the superscript ’$s1$’ signifies thresholds related to these six electrodes. If the capacitance falls below the threshold, indicating insufficient contact, the control system adjusts the finger’s position to re-establish “full contact”. The regulation of the two Dynamixel motors, which operate the two soft fingers via tendons, is defined as:
@@
Figure 3: Capacitance measurements from all eight capacitive electrodes as the robot maneuvers the SkinGrip around an individual’s limbs. (a) & (b) Capacitance values of six capacitive electrodes during various displacements of the tendon line as the manipulator wraps around the arm and lower leg, respectively. (c) & (d) Capacitance values of the eight capacitive electrodes as the SkinGrip approaches, wraps around, and maintains dynamic “full contact” with the arm or lower leg, respectively. We normalize all signals to the range of [0, 1].
Mobile Manipulator Control In this section, we discuss the capacitive servoing control scheme employed to maneuver the Stretch RE1 end effector along a human limb.
As mentioned above, we define a threshold $\bm{C}_{th}^{s1}\in\mathbb{R}^{6}$ to ensure “full contact” between the two soft fingers and the human skin. Additionally, we require a threshold ${C}_{th}^{s2}\in\mathbb{R}$ for the $7$-left inner electrode. This is crucial to detecting contact along the approach axis between the upper part of the soft finger and human skin. Leveraging the left and right inner electrodes (Fig. 2 (c)), the two-electrode capacitive sensor is capable of measuring roll rotation $\gamma$, Fig. 2 (b), on the limb’s surface. Note that our system does not infer translations along the $x$- and $y$-axes or rotations ($\alpha$ and $\beta$) around the $z$-axis and $y$-axis (Fig. 2 (b)). Instead, the robot translates its end effector at a constant velocity of 4 cm/s along the central axis of a human limb to ensure bathing task progress.
Next, we carried out tests to monitor the variations in capacitance values as the SkinGrip approaches a human limb and encloses the limb until “full contact” is established between the two soft fingers and the human skin. This process is illustrated in Fig. 3 (c) and (d). The end effector continues to approach the human limb with its fingertips open so long as the value of the $7$-left inner electrode (Fig. 2 (c)) falls below the threshold $C_{th}^{s2}$, indicating a lack of contact between the upper portion of the soft finger and the human skin. The control system adjusts the manipulator’s position $z_{t}$ in (2) to regulate the touch interaction with the skin.
When the value of the left inner sensor reaches the threshold $C_{th}^{s2}$, the two soft fingers commence a closure phase. During the closure phase the two fingers progressively close and surround the human limb for “full contact”. The closure phase ends when the sum of errors between the current and desired values of the six electrodes is minimized, indicating “full contact”. This is achieved by dynamically adjusting the motor to bend the fingers. The system then maintains this “full contact” dynamically, using a proportional controller (see 1) for continuous feedback and adjustment, and continuously adapting $z_{t}$ and $\gamma_{t}$ according with the P controller in (2). This ensures the fingers adapt to changes in limb position or size, maintaining contact by adjusting tendon tension based on feedback from all electrodes.

### v2

# SkinGrip: An Adaptive Soft Robotic Manipulator with Capacitive Sensing for Whole-Limb Bed Bathing Assistance

@@ Abstract
Robotics presents a promising opportunity for enhancing bathing assistance, potentially to alleviate labor shortages and reduce care costs, while offering consistent and gentle care for individuals with physical disabilities. However, ensuring flexible and efficient cleaning of the human body poses challenges as it involves direct physical contact between the human and the robot, and necessitates simple, safe, and effective control. In this paper, we introduce a soft, expandable robotic manipulator with embedded capacitive proximity sensing arrays, designed for safe and efficient bed bathing assistance. We conduct a thorough evaluation of our soft manipulator, comparing it with a baseline rigid end effector in a human study involving $12$ participants across $96$ bathing trails. Our soft manipulator achieves an an average cleaning effectiveness of $88.8\%$ on arms and $81.4\%$ on legs, far exceeding the performance of the baseline. Participant feedback further validates the manipulator’s ability to maintain safety, comfort, and thorough cleaning. Video demonstrations are available on our project website  [^1] https://sites.google.com/view/softbathing.
@@ I Introduction
Prior literature has introduced a number of robotic devices to assist a person in bathing their body. The majority of prior work introduces rigid flat bathing tools to contact and clean small surface regions of human limbs [7, 8, 9, 10]. However, these devices present certain challenges: the use of rigid bathing tools that come in contact with human skin may pose safety risks during interactions with individuals who require assistance, and achieving effective cleaning of a limb’s entire cylindrical surface often necessitates large and complex movements with a rigid end effector. Soft deformable devices present a promising alternative in assistive robotics, attributed to their lightweight, safe touch, and high adaptability [11, 12].
@@
In this work, we develop a robotic system to improve safety in human-robot interactions, specifically designed to aid in bed bathing tasks, as shown in Fig. 1. This system includes a novel soft robotic manipulator—SkinGrip, equipped with two tendon-driven soft fingers. The two fingers are uniquely structured to gently encircle human limbs. Additionally, we present a capacitive servoing control scheme. This approach empowers the SkinGrip to facilitate safe and continuous contact with human limbs, when used with a Stretch RE1 mobile manipulator. Finally, we conduct a robotic bathing human study experiment to evaluate our SkinGrip against a baseline end effector [7, 9], evaluating cleaning performance and participant perspectives, across participants with varying sized limbs. The study results demonstrate that the proposed manipulator and control strategy were able to efficiently clean the entire surface of a human limb with an average cleaning effectiveness of $88.8\%$ on arms and $81.4\%$ on legs, surpassing the baseline in both cleaning efficiency and user experience. In this work our contributions include:
@@
    We implement a control approach for the SkinGrip that ensures effective skin bathing performance and safe physical interaction. This method integrates multidimensional capacitive sensing to maintain dynamic, full contact with the skin.
@@ II-A Robot-Assisted Bed Bathing
Bed bathing is a crucial aspect of nursing care, and various robot-assisted approaches have been proposed to enhance this process. Zlatintsi and Dometios et al. [13, 14, 15, 16, 17, 18] developed a vision-based system with a soft arm for washing or wiping the back region of a person. King et al. [8] showcased an end effector designed for limb cleaning, operated via a point cloud interface for area selection. Erickson et al. [7, 9] utilized capacitive sensing to track human motion, enabling a mobile manipulator to move a rigid end effector along a human limb. Liu et al. [19] introduced a mesoscale wearable robot that cleans human limbs by navigating over them. Recently, Madan et al. [20] presented RABBIT, a unique robot-assisted bed bathing system that employs multimodal perception and dual compliance for a robotic manipulator to navigate a flat bathing tool along a human limb. Gu et al. [21] introduced a visuo-tactile Transformer-based imitation learning method (VTTB) for bathing assistance, employing multimodal sensing with a Baxter robot to track body contours and perform precise actions across different subjects.
In contrast to this body of prior literature, the proposed soft end effector design in this paper is able to make contact with and bath an entire limb circumference. To date, prior literature has leveraged a planar bathing instrument that is capable of making contact with only a small region of the skin at once [7, 20, 21, 10]. In addition, we introduce a capacitive proximity sensing approach for soft robot fingers that enable maintaining full-surface contact with the human body to improve cleaning effectiveness.
@@ II-B Soft Manipulators for Healthcare
Soft robotics has become increasingly prevalent in healthcare and biomedical fields, especially for rehabilitation and assistance. A comprehensive review [22] examines soft robotic hands designed for rehabilitation, revealing that $34\%$ of the $44$ unique devices analyzed are tendon-driven. Polygerinos et al. [23] developed an electromyography (EMG)-activated soft robotic glove to aid in grasping objects for ADLs. Cianchetti et al. [24] provide a comprehensive overview of soft robotics in biomedical applications, highlighting compliance, safety, and effectiveness in gentle interactions with human limbs, especially under high contact forces. Furthermore, advancements in soft robotics have extended to assistive devices for bathing. Zlatintsi et al. [15] introduced a fluid and tendon-operated soft robotic arm within the I-Support bathing system. Huang et al. [25] presented a soft pneumatic robotic tool for wiping and bathing that can adapt to slight body contours, but only makes contact with small regions of skin at a time.
In contrast to prior literature, we introduce a tendon-driven soft mechanism to achieve full limb circumference bed bathing assistance. Furthermore, we integrate capacitive proximity sensing within the soft finger architecture to sense distance from and contact with the human body. Capacitive proximity sensing enables the soft fingertips to dynamically adjust their curvature (through actuation) to ensure continuous contact and uniform pressure distribution between the finger and a human limb.
@@ II-C Capacitive Servoing and Control
Capacitive sensing is widely utilized in healthcare for applications ranging from classifying gait phases with shoe insole sensors [26] to capturing clothing state as a person is getting dressed using capacitive sensors embedded in the garment [27]. Similarly, Erickson et al. [7, 9] developed a capacitive servoing method using an array of capacitive sensors on a flat rigid tool for robotic bathing and dressing assistance. This approach enabled a robot end effector to traverse the contours of a human limb, but was only demonstrated to clean off the top surface of a limb. Building on these concepts, our work also employs capacitive proximity sensing, but focuses on defining control trajectories for a new manipulator designed with soft fingers that can safely encircle and clean the entire circumference of a limb, addressing the limitations of previous designs.
@@ III Design, Fabrication, and Capacitive Sensing
Figure 2: 3D model and prototype of the proposed SkinGrip, as well as the baseline end effector. (a) The SkinGrip is mounted on a Stretch RE1 robot for cleaning tasks. (b) 3D model of the SkinGrip. (c) Real-world SkinGrip, equipped with eight copper foils labeled as: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, $6$-right bottom, $7$-left inner, and $8$-right inner. (d) & (e) The SkinGrip equipped with a sponge and then wrapped with a washcloth. (f) 3D model of the baseline end effector. (g) Real-world baseline end effector, equipped with six copper foils labeled as: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, and $6$-right bottom. (h) & (i) The baseline end effector equipped with a sponge and then wrapped with a washcloth. For both the SkinGrip and the baseline end effector, the capacitive sensors are encased in a protective plastic film to prevent direct contact with the washcloth and human skin.
III-A Design, Sensing, and Actuation of the SkinGrip
Soft Gripper Design and Fabrication The proposed SkinGrip design is illustrated in Fig. 2(b)-(e). It comprises two soft fingers designed to wrap around a human limb. The size of the soft gripper was selected based on anthropometric data [28] to accommodate a wide range of adult limb diameters ($8$ cm to $16$ cm), ensuring it can effectively wrap around and clean the entire circumference of the limb. We optimized the width and thickness of the soft fingers to balance friction and contact area: a width that minimizes excessive friction while maintaining sufficient contact area, and a thickness that ensures flexibility without causing discomfort. Consequently, the final dimensions of the soft gripper are $160$ mm in diameter, $30$ mm in width, and $3$ mm in thickness. Each finger is tendon-actuated and powered by two Dynamixel XM430-W350-T motors. The motors bends the fingers by pulling tendons embedded within the soft material (Fig. 2(c)), causing them to curl inward and wrap around the limb when tensioned. The required bending curvature resembles a circular arc and adapts to the shape of the limb, ensuring full contact without causing discomfort or excessive pressure. Our approach is based on the theoretical model presented by [29], which provides the necessary equations of curvature.
@@ III-A Design, Sensing, and Actuation of the SkinGrip
Soft Finger Actuation and Capacitive Feedback In order to define a control strategy for the SkinGrip, we first quantify the relationship between capacitance measurements from the embedded electrodes and displacement of the fingers around a human limb. We have the two fingers of the SkinGrip wrap around an individual’s limbs, which remain stationary throughout the entire procedure. Fig. 3(a) and (b) depict the correlation between capacitance values and tendon line displacement as soft fingers wrap around a human arm and lower leg, respectively. The results show that capacitance values increase with the tendon’s displacement, indicating an increasing level of tightness between the soft finger and the limb. Full contact refers to the condition in which the inner surface of the soft finger is in complete contact with the human skin. Furthermore, as the servos continue to apply tension on the tendon, the soft finger compresses the human skin until it achieves tight contact, at which point the capacitance values stabilize. Notably, capacitance measurements from all sensors increase as the fingers continue to tighten around a human limb, and this trend is consistent for both arms and legs.
@@
The robot captures capacitance measurements from each electrode at a frequency of 10 Hz. At each time step, the robot will take control actions based on a windowed observation of the most recent capacitance measurements $\bm{C}^{s}_{t-4:t}\in\mathbb{R}^{8\times 5}$ from the capacitive sensor from the time step $t-4$ to the current time step $t$. Subsequently, we normalize all real-time capacitance values to the range $[0,1]$ using the formula $\bm{C}_{norm}^{s}=\frac{\sum\bm{C}^{s}_{t-4:t}/5-\bm{C}_{min}^{s}}{\bm{C}_{max}^{s}-\bm{C}_{min}^{s}}$ where the superscript $s$ represents the SkinGrip. Normalization of these values, determined by the maximum $\bm{C}_{max}^{s}$ and minimum $\bm{C}_{min}^{s}\in\mathbb{R}^{8}$ capacitance values, ensures consistent sensitivity across different conditions. We establish a threshold $\bm{C}_{th}^{s1}\in\mathbb{R}^{6}$ for six of the eight electrodes, specifically: $1$-left top, $2$-left middle, $3$-left bottom, $4$-right top, $5$-right middle, and $6$-right bottom (Fig. 2(c)), to ensure full contact between the soft fingers and the human skin. Here, the superscript ‘$s1$’ signifies thresholds related to these six electrodes. If the capacitance falls below the threshold, indicating insufficient contact, the control system adjusts the finger’s position to re-establish full contact. The regulation of the two Dynamixel motors, which operate the two soft fingers via tendons, is defined as:
@@
Figure 3: Capacitance measurements from all eight capacitive electrodes as the robot maneuvers the SkinGrip around an individual’s limbs. (a) & (b) Capacitance values of six capacitive electrodes during various displacements of the tendon line as the manipulator wraps around the arm and lower leg, respectively. (c) & (d) Capacitance values of the eight capacitive electrodes as the SkinGrip approaches, wraps around, and maintains dynamic full contact with the arm or lower leg, respectively. We normalize all signals to the range of [0, 1].
Baseline Rigid End Effector Design We contrast the SkinGrip to a baseline rigid flat end effector based on prior work [7, 9], shown in Fig. 2(f)-(i). This baseline end effector consists of six capacitive electrodes arranged in a $2\times 3$ grid, adhered to the bottom surface of the rigid tool, as shown in Fig. 2(g). These copper foils are connected to a Teensy LC microcontroller board. The flat tool is also covered with shape memory foam and a washcloth, shown in Fig. 2(h)-(i). This rigid flat tool was manufactured using a Raise3D E2 printer with PLA material. Finally, we mount the two end effectors to a Stretch RE1 mobile manipulator to perform body bathing along the human limbs, as shown in Fig. 2(a).
III-B End Effector Operation from Capacitive Sensing
Our goal is to instruct the Stretch RE1 to maneuver both the SkinGrip and the baseline end effector along a human limb for bathing assistance. The robot adjusts the end effector’s position and orientation using capacitance measurements from the integrated proximity sensors, enabling adaptive interaction with the human limb.
SkinGrip Operation As mentioned in Section III-A, we define a threshold $\bm{C}_{th}^{s1}\in\mathbb{R}^{6}$ to ensure full contact between the two soft fingers and the human skin. Additionally, we require a threshold ${C}_{th}^{s2}\in\mathbb{R}$ for the $7$-left inner electrode. This is crucial to detecting contact along the approach axis between the upper part of the soft finger and human skin. Leveraging the left and right inner electrodes (Fig. 2 (c)), the two-electrode capacitive sensor is capable of measuring roll rotation $\gamma$, Fig. 2 (b), on the limb’s surface. Note that our system does not infer translations along the $x$- and $y$-axes or rotations ($\alpha$ and $\beta$) around the $z$-axis and $y$-axis (Fig. 2 (b)). Instead, the robot translates its end effector at a constant velocity of 4 cm/s along the central axis of a human limb to ensure bathing task progress.